Southeast Airlines had pretax earnings of $80 million. Included in this amount is income from discontinued operations of $10 million. The company’s tax rate is 25%.

What is the amount of income tax expense that Southeast would report in its income statement for continuing operations? (Enter your answer in millions rounded to 2 decimal place (i.e., i.e., 5,500,000 should be entered as 5.50). Amount to be deducted should be indicated with a minus sign.)

Income from continuing operations before tax $70.00selected answer correct
Income tax expense 20.00selected answer incorrect
Income from continuing operations $90.00
Income from discontinued operations 2.50selected answer incorrect
Net income
Answer #1
Income from continuing operations before tax             70.00
Income tax expense             17.50
Income from continuing operations                 52.50
Income from discontinued operations               7.50
Net income             60.00

Explanation: 17.50 = 25% of 70

Thus income from continuing operations = 70-17.50 = 52.50

Income from discontinued operations = 10 - 25% of 10

= 7.50

Net income = 52.50 + 7.50 = 60